# New Perspectives in Advancing Graph Machine Learning

### v1

@@ 4 Experimental Results
We conduct experiments with an Equivariant Scattering-based GNN (ESc-GNN), illustrated in Figure 3, based on the vector-valued diffusion wavelets and scattering transform introduced in Section 3. We consider node-level and graph-level tasks on synthetic $3D$ point clouds randomly sampled from the surface of randomly generated ellipsoids. As baselines, we used ESc-GNN (non-equivariant), an ablated version of our model which treats each vector-valued signal $\mathbf{w}:V\rightarrow\mathbb{R}^{3}$ as three separate scalar-valued signals; a non-equivariant scattering based GNN, LEGS (Tong et al., 2020); two equivariant GNNs, EGNN (Satorras et al., 2021) and TFN (Thomas et al., 2018); and several standard message passing networks, GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), and GIN (Xu et al., 2019). (For a detailed description of our architecture and experimental setup, as well as a discussion of computational complexity, please see Appendices D and E.) For node features, we use the $3D$ coordinates of the points. Collectively, our experiments demonstrate the importance of rotational equivariance, with our ESc-GNN  achieving comparable performance to other equivariant GNNs with significantly fewer trainable parameters. [^5] Our code is available at https://github.com/dj408/esc-gnn
@@
We first consider a graph-level task, on a data set of $N_{G}=512$ graphs, where the goal is to predict the Euclidean diameter of each graph, defined by $\text{diam}(G_{j})=\max_{v,v^{\prime}\in V(G_{j})}\|v-v^{\prime}\|_{2}$. To construct these graphs, we define random ellipsoids $\mathcal{E}_{j}=\{(x,y,z)\in\mathbb{R}^{3}:x^{2}/a_{j}^{2}+y^{2}/b_{j}^{2}+z^{2}/c_{j}^{2}=1\}$, $1\leq j\leq N_{G}$, where $a_{j}$, $b_{j}$, and $c_{j}$ are i.i.d. randomly generated coefficients with each $a_{j}\sim\mathcal{N}(3,0.5)$ and $b_{j},c_{j}\sim\mathcal{N}(1,0.2)$, and, for convenience, we use $x,y$ and $z$ to denote the coordinates of a point in $3D$ space, i.e., $v=(x,y,z)$. We note that by construction, we will usually have $a_{j}^{2}>b^{2}_{j},c_{j}^{2}$ so these ellipsoids will typically be more “stretched out" along the $x$-axis than along the $y$ and $z$ axes.
@@
From each ellipsoid $\mathcal{E}_{j}$, we then sample $n=128$ points, $\{v^{(j)}_{i}\}_{i=1}^{n}$, and construct a $k$-NN graph with $k=5$. To demonstrate the utility of equivariant models, we rotate each ellipsoid in the test set (and only in the test set) by $90$ degrees so that they are most stretched out along the $y$ axis. Our results are shown in Table 1. We observe that while the non-equivariant models are able to perform well on the (non-rotated) validation set, they fail spectacularly on the test set. For example, the non-equivariant version of ESc-GNN  has an average validation mean square error (MSE) of $0.3126$, but average test MSE of $2.9181$. By contrast, the equivariant version of ESc-GNN, as well as the other equivariant networks, do not suffer from this limitation and achieve similar performance on validation and test sets. Additionally, we note that ESc-GNN  outperforms the equivariant baselines with only $59,779$ parameters compared to the $284,036$ parameters in EGNN and the $29,672,961$ parameters in TFN.
@@
We next consider a node-level task where the goal is to learn the value of a function $\mathbf{h}:\mathcal{E}\rightarrow\mathbb{R}^{3}$ defined on the underlying ellipsoid. At each vertex $v_{i}$, the direction of $\mathbf{h}(v_{i})$ is chosen to be the outward normal vector of $\mathcal{E}$. The magnitude is constructed via a randomly generated bandlimited function, generated using the first $K=16$ eigenfunctions of the Laplace-Beltrami operator. (See Appendix E.3 for details.) Overall, our results, shown in Table 2, tell a similar story to the graph-level results in Table 1, with the non-equivariant methods perform significantly worse on the test set than on the (non-rotated) validation set (although the drop in performance is not quite as extreme as before). The performance of ESc-GNNis comparable to the other equivariant GNNs, slightly better than EGNN and slightly worse than TFN. However, ESc-GNN  uses only $57,250$ parameters in comparison to $834,824$ for EGNN and $29,664,384$ for TFN. Therefore, these results indicate that ESc-GNN  is a fast, lightweight alternative to standard equivariant GNNs.
@@
Table 1: Results on the diameter prediction task (mean $\pm$ std. across 5-fold CV). Validation MSE is the (average) of the best score achieved during training for each fold. Best is bolded; second-best is underlined. ESc-GNN outperforms other equivariant methods with a fraction of the parameter count.
@@
Table 2: Node-level results (mean $\pm$ std. across five-fold CV). Validation MSE is the (average) of the best score achieved during training for each fold. Best test MSE score is bolded; second-best is underlined. ESc-GNN is the second best method to TFN, but with less than 0.2% of the parameters.
@@ 5 Conclusion
We have introduced a novel version of the geometric scattering transform for vector-valued signals. We have proved theoretical results demonstrating the rotational equivariance of our method and conducted experiments showing that it may be effectively incorporated into a GNN, which achieves comparable performance to baselines with a fraction of the parameter count.
@@ Appendix A Proofs of Main Theorems
where $c$ a universal constant.
We will also need the following lemma which provides a simplified expression for $\mathbf{Q}^{m}$
@@
Figure 1: Illustration of rotational equivariance of $\mathbf{Q}^{m}$, applied to a $2D$ vector field. (The field was generated with random uniform sampling of points $x_{i}\sim\mathcal{U}([-1,1]^{2})$, angles $\theta_{i}\sim\mathcal{U}[0,2\pi)$, and magnitudes $m_{i}\sim\mathcal{U}[0.4,1.0]$ such that $w_{i}=m_{i}[\cos\theta_{i},\;\sin\theta_{i}]$). The central vector (the magnitude of which was made to be the largest, for illustrative purposes) is colored blue. The top row shows $\mathbf{Q}^{m}$ applied to vectors in an unrotated vector field (where we first used the sample points $x_{i}$ to construct a $k$-NN graph, $k=3$). The bottom row shows the same system, rotated 90 degrees counter-clockwise, then diffused by $\overline{\mathbf{Q}}$. The black vectors highlight those involved in the next diffusion step. That is, the vectors from which the central vector will receive (indirect) diffusion messages: first, its (symmetric) $k$-nearest neighbor vectors, then neighbors of neighbors, and so on. After three diffusion steps, the top (unrotated) system is rotated 90 degrees counter-clockwise, like the bottom system was initially. (This panel has a light blue background). Notably, this panel is identical, (other than the background) to the panel immediately below it thereby demonstrating the equivariance of $\mathbf{Q}^{m}$, since diffusing and then rotating yields the same result as rotating and then diffusing.
@@
For any $m\geq 0$, and any vector-valued node feature $\mathbf{w}$ we have
@@
Now, fix $0\leq j\leq J$, and observe that we may write $\widetilde{\mathbf{\Psi}}_{j}=\mathbf{Q}^{t_{1}}-\mathbf{Q}^{t_{2}}$ where $t_{1}=0$, $t_{2}=1$ if $j=0$ and otherwiswe we have $t_{1}=2^{j-1},t_{2}=2^{j}$.
@@
Figure 2: Illustration of rotational equivariance of vector diffusion wavelets applied to a vector field (which is defined the same way as in Figure 1). Here, the top row shows vector diffusion wavelets constructed from $\mathbf{Q}$ applied to vectors in the unrotated vector field; the bottom row shows the same system, rotated 90 degrees counter-clockwise, then diffused using wavelets constructed from $\overline{\mathbf{Q}}$. The black vectors again highlight those from which the central vector will receive (indirect) messages in the next diffusion step. After three wavelet diffusion steps, the top (unrotated) system is rotated 90 degrees counter-clockwise, like the bottom system was initially (this panel has a light blue background). The fact that the diffused vector fields in the rightmost column are equivalent shows the rotational equivariance of the wavelets.
@@
We argue by induction. For the case case, $m=1$, we use the assumption that $\mathbf{R}$ commutes with rotations to see,
@@
Now, assume the result holds for some $m\geq 1$. Then by the inductive hypothesis, we have

### v2

@@ 4 Experimental Results
We evaluate VDW-GNNs across three progressively complex experimental settings: synthetic point clouds, real-world wind velocity measurements, and neural activity data. Our experiments include both node-level and graph-level targets; scalar-valued and vector-valued targets; and inductive classification, regression, and representation learning tasks. Collectively, these experiments demonstrate that VDWs provide a flexible framework for vector-valued learning on geometric data. A detailed discussion of their computational complexity can be found in Appendix D. [^4] Our code is available at https://github.com/dj408/vdw-gnns.
4.1 Synthetic ellipsoids
@@
We first consider node-level and graph-level tasks on synthetic 3D point clouds randomly sampled from the surface of randomly generated ellipsoids. For this task, we use a VDW-GNN model, which features vector- and scalar-track scattering modules with equivariant hidden feature mixing layers as shown in Figure 2. As baselines, we use a VDW-GNN (non-equivariant), an ablated version of our model which treats each vector-valued signal $\mathbf{w}:V\rightarrow\mathbb{R}^{3}$ as three separate scalar-valued signals; a non-equivariant scattering based GNN, LEGS (Tong et al., 2020); two equivariant GNNs, EGNN (Satorras et al., 2021) and TFN (Thomas et al., 2018); and several standard message passing networks, GCN (Kipf and Welling, 2017), GAT (Veličković et al., 2018), and GIN (Xu et al., 2019). For a detailed description of our architecture and experimental setup, see Appendix E.
For node features, we use the 3D coordinates of the points. Collectively, our experiments demonstrate the importance of rotational equivariance, with our VDW-GNN  achieving comparable performance to other equivariant GNNs with significantly fewer trainable parameters.
Figure 2: Model architecture variants for the ellipsoids experiments. Dotted arrows indicate that the scalar track’s hidden features can be used as input in the vector track’s mixing layer gate. Individual module descriptions can be found in Appendix E.1.
@@
We first consider a graph-level task, on a data set of $N_{G}=512$ graphs, where the goal is to predict the Euclidean diameter of each graph, $\text{diam}(G_{j})=\max_{v,v^{\prime}\in V(G_{j})}\|v-v^{\prime}\|_{2}$. To construct these graphs, we define random ellipsoids $\mathcal{E}_{j}=\{(x,y,z)\in\mathbb{R}^{3}:x^{2}/a_{j}^{2}+y^{2}/b_{j}^{2}+z^{2}/c_{j}^{2}=1\}$, $1\leq j\leq N_{G}$, where $a_{j}$, $b_{j}$, and $c_{j}$ are i.i.d. randomly generated coefficients with each $a_{j}\sim\mathcal{N}(3,0.5)$ and $b_{j},c_{j}\sim\mathcal{N}(1,0.2)$, and, for convenience, we use $x,y$ and $z$ to denote the coordinates of a point in 3D space, i.e., $v=(x,y,z)$. We note that by construction, we usually have $a_{j}^{2}>b^{2}_{j},c_{j}^{2}$ so these ellipsoids are typically more “stretched out" along the $x$-axis than along the $y$ and $z$ axes.
@@
From each ellipsoid $\mathcal{E}_{j}$, we then sample $n=128$ points, $\{v^{(j)}_{i}\}_{i=1}^{n}$, and construct a $k$-NN graph with $k=5$. To demonstrate the utility of equivariant models, we rotate each ellipsoid in the test set (and only in the test set) by $90$ degrees so that they are most stretched out along the $y$ axis. Our results are shown in Table 1. We observe that while the non-equivariant models are able to perform well on the (non-rotated) validation set, they fail spectacularly on the test set. For example, the non-equivariant version of VDW-GNN  has an average validation mean square error (MSE) of $0.3126$, but average test MSE of $2.9181$. By contrast, the equivariant version of VDW-GNN, as well as the other equivariant networks, do not suffer from this limitation and achieve similar performance on the validation and test sets. Additionally, we note that VDW-GNN  outperforms the equivariant baselines with only $59,779$ parameters compared to the $284,036$ parameters in EGNN and the $29,672,961$ parameters in TFN.
@@
We next consider a node-level task where the goal is to learn the value of a function $\mathbf{h}:\mathcal{E}\rightarrow\mathbb{R}^{3}$ defined on the underlying ellipsoid. At each vertex $v_{i}$, the direction of $\mathbf{h}(v_{i})$ is chosen to be the outward normal vector of $\mathcal{E}$. The magnitude is constructed via a randomly generated bandlimited function, generated using the first $K=16$ eigenfunctions of the Laplace-Beltrami operator. (See Appendix E.4 for details.) Overall, our results, shown in Table 2, tell a similar story to the graph-level results in Table 1, with the non-equivariant methods performing significantly worse on the test set than on the (non-rotated) validation set (although the drop in performance is not quite as extreme as before). The performance of VDW-GNN is comparable to the other equivariant GNNs, slightly better than EGNN, and slightly worse than TFN. However, VDW-GNN uses only $57,250$ parameters in comparison to $834,824$ for EGNN and $29,664,384$ for TFN. These results therefore indicate that VDW-GNN  is a fast, lightweight alternative to standard equivariant GNNs.
@@
Table 1: Results on the diameter prediction task (mean $\pm$ std. across 5-fold CV). Validation MSE is the (average) of the best score achieved during training for each fold. Best is bolded; second-best is underlined. VDW-GNN outperforms other equivariant methods with a fraction of the parameter count.
@@
Table 2: Node-level results (mean $\pm$ std. across five-fold CV). Validation MSE is the (average) of the best score achieved during training for each fold. Best test MSE score is bolded; second-best is underlined. VDW-GNN is the second best method to TFN, but with less than 0.2% of the parameters.
@@
4.2 Wind field reconstruction
We next consider a real-world wind data set that was previously considered by Battiloro et al. (2024), who introduced a tangent-bundle neural network defined in terms of the connection Laplacian. This data set consists of the earth’s mean 10-meter wind field measurements for a single day (January 1, 2016), retrieved from the (NOAA Physical Sciences Laboratory, 2025). Each sample in the data set has zonal (latitude-aligned) and meridional (longitude-aligned) vector components, as well as the corresponding latitude/longitude coordinates for the location of the measurement. Similar to Battiloro et al. (2024), in our experiment, we aim to reconstruct masked wind velocity measurements. Practically, this task can interpreted as recovering lost signals in a sensor network degradation scenario (e.g., random remote wind sensor failures across the globe).
We first subsample the data set to 2000 wind measurements. Then, we split those 2000 measurements, 70/30, into observed and masked sets, and further split the masked set 10/10/10 into masked-training, masked-validation, and masked-test sets. For all points in the masked set, we replace their wind vector features with the component-wise mean vector of the observed set, and set the reconstruction target as the original wind vector. Here, we wish to demonstrate the utility of our method for processing features taking values in the tangent plane of the 2D sphere lying in 3D space. Accordingly, we embed measurement locations on the unit sphere by converting latitude/longitude to 3D Earth-centered, Earth-fixed (ECEF) Cartesian coordinates, and normalize by the radius of the Earth (assuming sphericality). We also lift the 2D wind vector features into the same 3D global coordinate system (details in Appendix F).
We then construct a modified $k$-NN graph ($k=3$) using Euclidean distance in 3D in which (i) observed nodes are connected to their $k$ nearest observed neighbors with symmetric (undirected) edges, and (ii) masked nodes receive only directed incoming edges from their $k$ nearest observed nodes, observed $\to$ masked, (i.e., we take $\mathbf{D}$ to be the out-degree matrix when computing $\mathbf{P}$). This modification prevents non-informative messages masked nodes from averaging with the observed set. Edge weights are initialized as inverse distance, $w_{ij}=1/d_{ij}$, and then normalized by target node so that each node’s incoming weights sum to $0.5$.
We train all models with the mean squared error (MSE) loss function, using the MSE of the masked-validation set to enforce early stopping and prevent overfitting. We evaluate the models both on their MSE on the masked-test set wind measurements, and (akin to the ellipsoids experiments in Section 4.1) we compute their MSE after a random 3D rotation is applied to the test set. In the latter case, we use rejection sampling of randomly generated candidate 3D rotation matrices, until a matrix with a rotation angle between 90 and 160 degrees is found.
We again test LEGS, GCN, GAT, GIN, EGNN, and TFN as baselines, as well as the DD-TNN (‘Domain-Discretized Tangent Bundle Neural Network’) from Battiloro et al. (2024) which also considered wind field reconstruction on this data set. We note that somewhat similar to our method, DD-TNN uses local SVD to define matrices $\mathcal{O}_{i,j}$ which are then used to construct an approximation of the connection Laplacian, also referred to as a sheaf Laplacian. DD-TNN then uses this sheaf Laplacian to define a spectral convolutional network. Note that in order to implement EGNN and TFN for this task, we decouple (i) measurement points’ global positions, from (ii) these points’ wind vectors, such that message passing occurs between ‘position on earth’ neighbors, but messages update in ‘wind vector space’. We repeat our experiment five times with unique sampling and rotation seeds. More experimental details can be found in Appendix F. We report aggregated test set results in Table 3.
For this task, we use a simple VDW-GNN model in which we apply a single layer of vector diffusion wavelets and then, for each node, concatenate the 3D coordinates of its three closest neighbors. We then feed this transformed representation into a three-layer MLP. (Preliminary experiments indicated that this outperformed more complex models on this data set.)
Table 3: Wind field reconstruction task results on the masked-test set, (mean $\pm$ std.) across five repetitions. Best test MSE score is bolded; second-best is underlined. VDW-GNN again achieves best or second-best results among comparable models with far fewer parameters.
Model MSE $\downarrow$ Rotated MSE $\downarrow$ Sec. per epoch Best epoch Parameter count
[…]
LEGS 11.2551 ± 2.3892 27.5993 ± 9.2066 0.0080 ± 0.0003 258 ± 132 21,319
[…]
GAT 3.2784 ± 0.2189 3.8239 ± 0.1370 0.0053 ± 0.0039 54 ± 45 34,179
GIN 3.1754 ± 0.1940 3.5410 ± 0.3372 0.0045 ± 0.0030 20 ± 4 50,435
$D$-TNN 12.7592 ± 0.9288 14.8010 ± 1.8416 0.0114 ± 0.0003 118 ± 110 17,374
[…]
TFN (2-layer) 3.9898 ± 0.1895 3.9898 ± 0.1895 0.0128 ± 0.0034 88 ± 65 729,184
Our model achieves the best test set error and second-best rotated-test set error, with far fewer parameters than its competitors, EGNN and GCN. It substantially outperforms DD-TNN, which also was designed to process tangent-bundle valued features. Our model also achieves the fastest per-epoch training time, with the important caveat that the diffusion operator $\mathbf{Q}$ is constructed once and cached before training, which takes approximately 1.1 seconds on the wind data set. [^5] Our model’s per-epoch time cost could be reduced even further in the single-layer case by precomputing the wavelet transform step.
4.3 Multi-channel neural recordings
We next consider a neural activity data set from experiments by Kaufman et al. (2016) and processed by Gosztolai (2023), used in three recent works that developed representation learning models for neural data: MAnifold Representation Basis LEarning (MARBLE) (Gosztolai et al., 2025), Consistent EmBeddings of high-dimensional Recordings using Auxiliary variables (CEBRA) (Schneider et al., 2023), and Latent Factor Analysis for Dynamical Systems (LFADS) (Pandarinath et al., 2018). (More details in Appendix G.)
The data set consists of 24-channel neural activity time-series data recorded as a trained macaque (a species of monkey native to Asia and North Africa) physically moved their hand on a screen towards one of seven targets (‘conditions’) in different 2D directions from a central position in order to obtain a reward. It contains data from 44 separate experimental days, where each day has between 189 and 607 trials (mean: 333; total: 14,660). Formally, the $j$-th trial from day $i$ corresponds to a time series of length $T$ taking values in $\mathbb{R}^{24}$, i.e., $\mathbf{x}^{(i,j)}:[T]\rightarrow\mathbb{R}^{24}$. The modeling task is to learn low-dimensional embeddings of the per-timepoint neural signals within trials and days. Embedding quality is assessed by using the embeddings as input to a classifier.
We use the version of the data set provided at Gosztolai (2023), and we generally follow the same preprocessing methods as Gosztolai et al. (2025), except we do not apply PCA to the processed 24-dimensional features (for any model). Our setup also differs from previous treatments of this data in that we conduct our experiments in an inductive (rather than transductive) setting. This represents a more challenging task, and ensures that model embeddings reflect a generalization of the representation function to unseen trials, rather than memorization of the training graph.
As in Gosztolai et al. (2025), we construct a continuous $k$-nearest neighbors (CkNN) graph (Berry and Sauer, 2019), whose nodes are per-timepoint samples in a multi-trial, within-day neural state space and the edge set is constructed to have a density-adjusted number of neighbors. To construct vector-valued node features, we again follow the lead of Gosztolai et al. (2025) by numerically differentiating the neural activity at each node to obtain 24-dimensional “neural velocities." Our model applies vector diffusion wavelets to these velocities, concatenates wavelets’ outputs, feeds them into a four-layer MLP, and trains with a supervised contrastive loss (Khosla et al., 2020) function, combined with a custom, hard-negatives-focused sampling module.
Performance is assessed using the same probing methodology in Gosztolai et al. (2025): after representation learning a three-dimensional embedding of timepoints (the most challenging setting used in Gosztolai et al. (2025)’s experiments), we flatten and concatenate timepoint embeddings per trial, train a support vector machine (SVM) on these aggregated embeddings, and evaluate classification accuracy on held-out trial aggregated embeddings. We report results aggregated across all 44 days (performing five-fold cross-validation on separate models per day). As shown in Table 4, our VDW-GNN model achieves an overall classification accuracy of $66\%$, slightly out-performing MARBLE, which achieves 65% accuracy. However, across experimental days, these models’ results distributions are not significantly different under a two-sided Wilcoxon test (see Figure 5, in Appendix G). Nonetheless, our model both trains much faster and uses far fewer parameters than MARBLE, and significantly outperforms CEBRA and LFADS. For further details on the data set, our experimental setup, and each of the models, see Appendix G.
Table 4: Summary results for the neural data experiment. Classifier accuracy: the mean (across 44 days) of mean SVM classifier accuracies (across five-fold CV within days), $\pm$ standard deviation (of daily CV means). Best is bolded; second-best is underlined.
Model Classifier accuracy $\uparrow$ Sec. per epoch Best epoch Parameter count
VDW-GNN 0.66 ± 0.14 0.23 ± 0.05 73 ± 172 164,868
[…]
CEBRA 0.57 ± 0.09 0.40 ± 0.04 685 ± 3040 11,171
LFADS 0.35 ± 0.08 0.44 ± 0.11 5 ± 34 1,068,800
@@ 5 Conclusion
We have introduced a novel version of the diffusion wavelets for vector-valued signals, and proved theoretical results demonstrating their frame properties and rotational equivariance. Empirically, we have shown that these wavelets may be incorporated into geometric graph neural networks and effectively applied to synthetic point cloud data sets, as well as real-world data sets derived from wind field measurements or neural activity data.
Acknowledgements
DJ, MP, and SK were funded in part by NSF-DMS-2327211; AS by NSF-DMS-2136090; MP by NSF-OIA-2242769; DN by NSF-DMS-2408912. SK was also funded in part by NSF Career Grant 2047856.
@@ Appendix A Proofs of Main Theorems
In this section, we will provide the proofs of Theorems 3.1, 3.2, and 3.3. Additionally we recall that, as noted in Section 3, in the scalar-valued signal case, Tong et al. (2024) showed that the scattering transform could be extended to include non-dyadic diffusion wavelets of the form $\mathbf{P}^{t_{j}}-\mathbf{P}^{t_{j+1}}$ where $0=t_{0}<t_{1}<t_{2}\ldots$ is a generic sequence of increasing integers and that vector diffusion wavelets may be readily extended to non-dyadic scales. We note that it is straightforward to extend all of our theoretical analysis to wavelets and scattering transforms constructed in this way.
@@
where $c>0$ is a universal constant.
We also need the following lemma which provides a simplified expression for $\mathbf{Q}^{m}$.
@@
Figure 3: Illustration of rotational equivariance of $\mathbf{Q}^{m}$, applied to a 2D vector field. The vector field was generated with random uniform sampling of points $x_{i}\sim\mathcal{U}([-1,1]^{2})$, angles $\theta_{i}\sim\mathcal{U}[0,2\pi)$, and magnitudes $m_{i}\sim\mathcal{U}[0.4,1.0]$ such that $w_{i}=m_{i}[\cos\theta_{i},\;\sin\theta_{i}]$. The central vector (the magnitude of which was made to be the largest for illustrative purposes) is colored blue. The top row shows $\mathbf{Q}^{m}$ applied to vectors in an unrotated vector field where we first used the sample points $x_{i}$ to construct a $k$-NN graph, $k=3$. The bottom row shows the same system, rotated 90 degrees counter-clockwise, then diffused by $\overline{\mathbf{Q}}$. The black vectors highlight those involved in the next diffusion step, specifically, the vectors from which the central vector will receive (indirect) diffusion messages: first, its (symmetric) $k$-nearest neighbor vectors, then neighbors of neighbors, and so on. After three diffusion steps, the top (unrotated) system is rotated 90 degrees counter-clockwise, like the bottom system was initially. This is shown in the top-right panel with a tinted background. Notably, this panel is identical to the panel immediately below it, thereby demonstrating the equivariance of $\mathbf{Q}^{m}$, since diffusing and then rotating yields the same result as rotating and then diffusing.
@@
For any $m\geq 0$, and any vector-valued node feature $\mathbf{w}\in\mathbb{R}^{nD}$ we have
@@
Now, fix $0\leq j\leq J$, and observe that we may write $\widetilde{\mathbf{\Psi}}_{j}=\mathbf{Q}^{t_{1}}-\mathbf{Q}^{t_{2}}$ where $t_{1}=0$, $t_{2}=1$ if $j=0$ and otherwise we have $t_{1}=2^{j-1},t_{2}=2^{j}$.
@@
Figure 4: Illustration of rotational equivariance of vector diffusion wavelets applied to a vector field defined as in Figure 3). Here, the top row shows vector diffusion wavelets constructed from $\mathbf{Q}$ applied to vectors in the unrotated vector field; the bottom row shows the same system, rotated 90 degrees counter-clockwise, then diffused using wavelets constructed from $\overline{\mathbf{Q}}$. The black vectors again highlight those from which the central vector will receive (indirect) messages in the next diffusion step. After three wavelet diffusion steps, the top (unrotated) system is rotated 90 degrees counter-clockwise, like the bottom system was initially, as shown in the top-right panel with the tinted background. The fact that the diffused vector fields in the rightmost column are equivalent shows the rotational equivariance of the wavelets.
@@
Before proving Theorem 3.3 we briefly discuss the assumption that $\sigma$ commutes with rotations. To understand why this condition is necessary, observe that Theorem 3.2 implies that
Therefore, Theorem 3.3 will not hold without this assumption. We also note that we may readily construct activations $\sigma$ satisfying this assumption by defining $\sigma(\mathbf{w})(v)=\sigma_{\text{radial}}(\|\mathbf{w}(v)\|_{2})\frac{\mathbf{w}(v)}{\|\mathbf{w}(v)\|_{2}}$ when $\mathbf{w}(v)\neq 0$, and $\sigma(\mathbf{w})(v)=0$ when $\mathbf{w}(v)=0$, where $\sigma_{\text{radial}}$ is any real-valued function.
@@
We argue by induction. For the case case, $m=1$, we use the assumption that $\mathbf{R}$ commutes with rotations to see that
@@
Now assume the result holds for some $m\geq 1$. Then by the inductive hypothesis, we have